OpenStudy (anonymous):

Simplify: e^(i(1-n)pi)

OpenStudy (anonymous):

\[e ^{i(1-n)\pi}\] and \[e ^{i(1+n)\pi}\]

OpenStudy (anonymous):

I know that \[e ^{i \pi}=(-1)^{n}\]

OpenStudy (anonymous):

and what about \[e ^{-i n \pi}=(-1)^{n}?\]

OpenStudy (mr.math):

\(e^{in\pi}=(-1)^n\) for \(n=1,2,3,..\).

OpenStudy (mr.math):

You can get from that \(e^{i\pi}=-1\).

OpenStudy (mr.math):

I wrote that above.

OpenStudy (anonymous):

Read his response properly, what you said about e^i*Pi is not correct.

OpenStudy (anonymous):

oops, but i have this: \[e ^{(1-n)\pi}=e ^{i \pi}e ^{-i n \pi}\] \[e ^{i \pi}e ^{-i n \pi}=(-1)^{n}(-1)^{n}=(-1)^{2n}=1???\]

OpenStudy (anonymous):

i forgot the "i"

OpenStudy (mr.math):

This is not true!

OpenStudy (mr.math):

\(e^{i(1-n)\pi}=e^{i\pi}e^{-in\pi}=-1(-1)^n=(-1)^{n+1}\), where \(n=1,2,3,..\).

OpenStudy (anonymous):

\[e^{i*\pi} = -1 , jessica\] that was your problem, you were thinking it was \[(-1)^n\] as well

OpenStudy (mr.math):

In other words you will have -1 for n even and 1 for n odd.

OpenStudy (anonymous):

owhhh sh*t! e^(i*pi) is -1 =.= and e^(i*n*pi)=(-1)^n

OpenStudy (anonymous):

stupid me!

OpenStudy (mr.math):

Haha finally :P

OpenStudy (anonymous):

thank you guys <3

OpenStudy (mr.math):

You're welcome.

OpenStudy (anonymous):

No problem =)